OpenStudy (anonymous):

how to solve radical 28 plus radical 63

OpenStudy (anonymous):

\[\sqrt{28}+ \sqrt{63}\]

OpenStudy (zepp):

You have to simplify both radical expression first ;)

OpenStudy (saifoo.khan):

What he said^

OpenStudy (anonymous):

haha thanks

hero (hero):

Hint: Since 28 = 4 x 7 and 63 = 9 x 7, then we can write the above as : \[\sqrt{ 4 \times 7} + \sqrt{9 \times 7}\] Also since \[\sqrt{a \times b} = \sqrt{a} \times \sqrt{b}\] we can re-write the above as \[\sqrt{4} \times \sqrt{7} + \sqrt{9} \times \sqrt{7}\]

hero (hero):

I find it amazing that just two seconds ago, no one was even here helping the poor young fellow, then out of the blue, as soon as I show up zepp and saifoo show up.

OpenStudy (zepp):

Saifoo and I were here since the problem was posted D:

hero (hero):

When I came, no one was here and no one had yet to post a solution. I don't doubt you two were here first, but why did you wait until I started typing a solution to type anything at all?

OpenStudy (anonymous):

My Hero! LOL

OpenStudy (saifoo.khan):

We should make a film on this.

hero (hero):

Saifoo, I don't want to hear your jokes right now.

OpenStudy (saifoo.khan):

Dont hear.. just read.

hero (hero):

I don't want to read them either.

OpenStudy (saifoo.khan):

but you said u dont want to hear them.

OpenStudy (anonymous):

LOL I am enjoying READING this

OpenStudy (saifoo.khan):

Everyone does. ;D

OpenStudy (anonymous):

lol

hero (hero):

I would say, "saifoo the clown likes to play around", but that would be a bad joke.

hero (hero):

So hostile.

OpenStudy (anonymous):

still enjoying :)

hero (hero):

Only in your head

OpenStudy (zepp):